How do I solve for x in this equation?
x + (x * 5%) = 3
More to the point, the problem definition is this:
How do I get back to the base number when the result was computed by adding the base number to a percentage times the base number (ie. I know the result and the percentage)?

Answer by mananth(16946)

x+(x*1/20)=3
20x+x / 20 =3
21x=60
x=60/21
..
2.857=x